# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

### v1

@@ 1.1. Positioning within the literature
Finally, a clarification is in order: the literature has adopted diverse definitions of the terms novelty, innovation, and invention, often depending on the degree of novelty introduced or the social and economic impact of the innovation. In physics and mathematics, innovation processes are typically modeled as sequences of observed elements. An item is labeled a novelty or innovation if it appears for the first time in a sequence of observations; otherwise, it is classified as a repetition of a specific element observed in the past. A classic example is provided by the analysis of word sequences in texts, as explored by Tria et al. (2014), where a novelty is defined as the first occurrence of a word in the sequence, and subsequent appearances of the same word are treated as repetitions. In this framework, suitable models are the Poisson-Dirichlet model (Pitman and Yor 1997) and the urn with triggering (Tria et al. 2014) for a single process and their corresponding versions with interaction (Fortini et al. 2018, Iacopini et al. 2020, Aletti et al. 2023a, 2025).
In contrast, in economics, innovation is frequently synonymous of patented invention. Here, a patent is, by definition, a novel item—never previously observed—but the emphasis is on whether the patent constitutes a successful innovation. Economic studies thus often model the probability of generating a significant innovation—what we refer to in this work as a “success”. Accordingly, in this paper, we use the terms innovation, novelty, and invention interchangeably, focusing primarily on whether a patented item meets the threshold of success (with respect to some measure), regardless of whether it constitutes a breakthrough or an incremental improvement. Moreover, in this framework, the natural class of models are the interacting Bernoulli processes, which our model belong to.
To the best of our knowledge, there are only few papers in the scientific literature regarding interacting (or interrelated) Bernoulli processes. The first one is Antelman (1972), which tries to develop a Bayesian method for making statistical inference on the parameters of two Bernoulli processes when the a-priori joint distribution that generates them exhibits dependence. More precisely, a family of closed-under sampling prior distributions, called the Dirichlet-beta family are introduced. However, these distributions result intractable and several approximation are necessarily proposed. Moreover, this study does not cover the case of more than two Bernoulli processes, as in our case. Another paper is Katselis et al. (2019), where the parameters of the Bernoulli distributions at time $t+1$ depend only on the status of the processes at the present time $t$ so that a multi-lag dependence is not allowed. Finally, in Pandit et al. (2019) the provided model relates the parameter of the Bernoulli distribution of a certain process at a given time-step to the history of all the processes over $L$ time lags by means of interaction parameters $\theta_{h,j,\ell}$, that capture how much process $j$ at time lag $\ell$ affects process $h$. Hence, this model and ours share the dependence of the parameter of the Bernoulli distribution for the future observations on the present and past observations of all the processes. However, important differences are present. First of all, in Pandit et al. (2019) the dependence on the past is up to a fixed number $L$ of lags, while in our model the dependence is on all the past observations. Moreover, in Pandit et al. (2019) the parameters $\theta_{h,j,\ell}$ do not depend on the time-step $t$, while in our model the corresponding quantities essentially are $\gamma_{j,h}/(c+t)$, that decreases along time and let us to get, for each process, a vanishing success probability and a power-law behavior of the number of successes, that we observe in the patent data set. Finally, the paper Pandit et al. (2019) focuses on a maximum likelihood estimation procedure for the model parameters $\theta_{h,j,\ell}\in{\mathbb{R}}^{N\times N\times L}$, that, since the high number of parameters, needs some sparsity condition on $\Theta=\{\theta_{h,j,\ell}\}$; while we provide a deep theoretical study of the model obtaining asymptotic results that can be verified on real data.
It is also worthwhile to stress that the theoretical results for our model follow from new general theoretical results, that, among other things, have the merit to provide a general mathematical framework where we can also include the results in the above quoted papers Tria et al. (2014), Aletti et al. (2023b) regarding the number of distinct observed novelties.
1.2. Structure of the paper
@@
The remainder of the paper is organized as follows. In Section 2, we introduce our primary novelty measure based on a category-level variant of the forward citation index, and define what constitutes a “success” in each technological field. Section 3 presents the proposed mathematical model of interacting Bernoulli processes, offering a theoretical framework to capture the dynamics of innovation across categories, and states the main asymptotic results. All the theoretical results are analytically proven in the appendix (supplementary material). Moreover, Section 3 highlights the implications of the model through a simulation exercise, related to the propagation effects of exogenous shocks. In Section 4, we validate the theoretical findings using a comprehensive dataset of patents from the GLOBAL PATSTAT database, and we estimate the structural parameters of the interaction matrix. Statistical inference on the interaction intensity under a mean-field assumption is also performed. Section 5 concludes.
2. Category forward citation index and definition of “success”
@@
As previously noted in Subsection 1.1, in economics the term “innovation” often refers to “patented invention”. Since patents are novel by definition, our focus lies on whether they represent a successful innovation. Thus, we use the terms innovation, novelty, and invention interchangeably, emphasizing their success rather than their degree of novelty.
@@
The number of citations a given patent receives (more briefly, forward citations) reflects the technological importance of the patent for the development of subsequent technologies (see Squicciarini et al. (2013) and reference therein). In other words, forward citations to a certain patent represent the range of later inventions that have benefited from that patent. Therefore, we have decided to use a variant of the forward citation index in Squicciarini et al. (2013) for detecting the patents that could be defined a “success” in one or more categories. However, it is worthwhile to note that our approach is not specific for a certain classification in categories or for the chosen index and so it could be also applied with a different index aimed to evaluate the importance of a patent. [^6] A review of the novelty indices present in the literature and potentially applicable to our context is presented in Appendix A. There we also motivate our preference for the index of Squicciarini et al. (2013) over other measures of patent novelty.
@@
$CIT_{n,h}$= number of patents with publication date in the period $[d_{n},d_{n}+365\times T]$ and belonging to category $h$ and citing $n$.
@@
Note that the number of forward citations a patent gets can differ greatly depending on the publication year and the category it comes from. This is why we decided to share the approach adopted in Squicciarini et al. (2013) and normalize this number with respect to the cohort of the patent considered. In other words, the index we use is a disaggregation of the forward citation index, say $I_{S}(n)$, in Squicciarini et al. (2013): indeed, we have
@@
Remark 2.1 (Invariance property of the index).
@@
According to the above definition, if we denote by $X_{n,h}$ the random variable that is equal to $1$ if patent $n$ is a success for category $h$ and equal to 00 otherwise, we obtain a finite system of Bernoulli processes, that is $\{X^{(h)}=(X_{n,h})_{n}\,:\,h\in{\mathcal{H}}\}$. These processes are not independent, but they interact. This interaction is crucial to consider, as numerous studies have highlighted that –within innovation ecosystems– innovation processes do not evolve in isolation but often influence one another through knowledge spillovers, recombination, and co-dependence, shaping both innovation outcomes and strategic investment decisions across domains (Adner and Feiler 2019, Bondarev and Krysiak 2021, Kim and Magee 2017).
@@
3. Mathematical model and theoretical results: Interacting Bernoulli processes
@@
(3.1) $P_{t,h}$ $=P(X_{t+1,h}=1\,|\,\text{past until time-step $t$})$
@@
where $\theta_{h}>0$, $c\geq\theta_{h}$ are parameters that tune the initial condition in the model dynamics (for details, see (B.1) in Appendix B) and $S_{t,j}=\sum_{n=1}^{t}X_{n,j}$ is the number of successes until time-step $t$ for process $j$. Moreover, the parameters $\gamma_{j,h}$ play the fundamental role in the dynamics of the system (see again (B.1)). We refer to the matrix $\Gamma=(\gamma_{j,h})_{j,h}$ as the interaction matrix and we assume the following conditions:
@@
[…]
@@
    $\Gamma$ is irreducible, that is the graph with the Bernoulli processes as nodes and with $\Gamma$ as the (weighted) adjacency matrix is strongly connected. (In Appendix C we will discuss the case of a reducible matrix $\Gamma$ [^7] This extension may be particularly relevant in contexts where the USPTO –or other patent offices—modify the structure of technological categories through reclassification (Lafond and Kim 2019, Chae and Gim 2019) thus changing the landscape of interaction among the latter.)
@@
In our context, given $t$ patents, the probability $P_{t,h}$ that the future patent $t+1$ will be a success in category $h$ can have an increasing dependence on the number of past successes in any category (self- and cross-category reinforcement) with the parameter $\gamma_{j,h}$ ruling the strength of this (direct) dependence and the assumed irreducibility of the matrix $\Gamma$ generates a global interaction among all the categories (according to the literature on knowledge spillovers among patents Jaffe et al. (2000), Thompson and Fox-Kean (2005)).
@@
  • in the case $\gamma^{*}<1$ (that is when ${\boldsymbol{1}}^{\top}\Gamma\neq{\boldsymbol{1}}^{\top}$) the probability of observing a success in process $h$ converges almost surely toward zero at the same rate $1/t^{(1-\gamma^{*})}$ for every process;
  • in the case $\gamma^{*}=1$ (that is when ${\boldsymbol{1}}^{\top}\Gamma={\boldsymbol{1}}^{\top}$) the probability of observing a success in process $h$ converges almost surely to a finite strictly positive random limit, which is the same for each process (note that in this case ${\mathbf{u}}$ is, up to a multiplicative non-zero constant, equal to the vector ${\boldsymbol{1}}$); in other words, at the steady state, the probability of having a success in a process is the same for each process of the system.
@@
Moreover, we recall that, according to the Perron-Frobenius theory, the components of the vector ${\mathbf{u}}$ are all different from zero and with the same sign and, in graph theory, they correspond to the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$). The eigenvector centrality is a measure of the importance of a node in a graph with respect to its out-links (if it is computed for the adjacency matrix) or its in-links (if it is computed for the transpose of the adjacency matrix). A high eigenvector centrality score means that the node points to, or respectively is pointed to by, many nodes with high scores. In the above theorems ${\mathbf{u}}$ is the vector of the relative eigenvector centrality scores with respect to $\Gamma^{\top}$ and, hence, by (3.1), a high value of $u_{h}$ means that the probability of having a success for process $h$ depends on the number of past successes observed in many processes who themselves have high scores.
@@
In short, a balanced innovation ecosystem ($\gamma^{*}=1$) supports ongoing, shared uncertainty and correlated success across technologies, which characterizes connected innovation ecosystems (Jacobides et al. 2024). In contrast, a not-balanced one ($\gamma^{*}<1$) leads to asymmetric, decoupled innovation dynamics over time. In other words, in well-connected systems, sustained uncertainty and interdependence across technological domains facilitate diverse learning paths and capability recombination. This interdependence fosters mutual reinforcement across domains, enabling firms to accumulate diverse, complementary capabilities. Conversely, in fragmented systems, limited spillovers may rigidify specialization patterns, potentially turning capabilities into rigidities over time.
@@
It is interesting to note that, in the innovation framework, Theorem 3.3 highlights how the long-run proportions of successful patents in each category reflect the connectedness and influence of that category within the broader innovation network. While this relationship is formally established in the present work, related insights have been suggested in the literature without formal proof. For example, Pichler et al. (2020) show that the innovation rate of a technological domain is influenced by the innovation rates of the domains it depends on. Similarly, Sampat and Ziedonis (2005), Zhang et al. (2025) emphasize the importance of the quality of a patent’s prior connections as well as their technological domain. In an effort to further deepen this latter aspect, we developed Theorem 3.4. As explained in the following, this theorem examines not only how many successful patents there are in each target category $h$, but also traces the source category $k$ to which each successful patent originally belonged. Namely, in our particular context, we can enrich the model assuming that the number $S_{t,k,h}$ of successes until time-step $t$ in category (process) $h$ coming from belonging category $k$ is of the form $S_{t,k,h}=\sum_{n=1}^{t}X_{n,h}Y_{n,k}$, where $Y_{n,k}$ takes value $1$ if the category of patent $n$ is $k$ and 00 otherwise and
@@
    ${\mathbf{Y}}_{n}=(Y_{n,k})^{\top}_{k\in{\mathcal{H}}}$ is independent of ${\mathbf{X}}_{n}=(X_{n,h})^{\top}_{h\in{\mathcal{H}}}$ and of all the past until time-step $n-1$ with $P(Y_{n,k}=1)=\pi_{k}\in(0,1)$ (where $\sum_{k\in{\mathcal{H}}}\pi_{k}=1$).
@@
To summarize, Theorem 3.4 provides a formal representation of how the structure of innovation inputs shapes the distribution of successful outcomes across technological domains. Specifically, it shows that, asymptotically, the share of successful innovations in a target category mirrors the origin distribution of patent activity. This result offers entrepreneurs a valuable mechanism for understanding how ecosystems embed and propagate specialization trajectories, which is a critical insight, given that the transformation of skills into core capabilities depends fundamentally on how technological domains interact and reinforce one another over time (Teece 2009, Jacobides et al. 2018, Kim and Magee 2017).
3.1. Model implication: simulations with a shock
Figure 1. Linear behavior of the successes counter $S_{t,1}$ (red line) and $S_{t,2}$ (blue line) along $t$, in $\log_{10}-\log_{10}$ scale, in two different scenarios with a positive shock on process $1$. The two scenarios are distinguished by the type of interaction, i.e. by the form of the interaction matrix $\Gamma$ used in the simulations. The shock on process $1$ occurs at time $t_{shock}=10^{4}$ and the system is observed until time-step $t=10^{7}$. We can observe the sudden rise of $S_{t,1}$ (red line) caused by the shock and how the positive effect of the shock on process $1$ propagates to process $2$ (consequent progressive rise of the blue line). We can also note that the slope of the two lines and the distance between their intercepts are the same before and after the shock, because the interaction matrix $\Gamma$ does not change: indeed, the slope is equal to $\gamma^{*}$ and the distance between the intercepts corresponds to the quantity $\log_{10}(u_{1}/u_{2})=\log_{10}(u_{1})-\log_{10}(u_{2})$.
In order to highlight what is the implication of the fact that a system follows the provided model, we here show how the effect of a shock on a process propagates to the other ones by means of the interaction in their dynamics. A vivid example of this phenomenon is the surge in pharmaceutical innovation during the COVID-19 pandemic, which triggered subsequent innovations in other fields such as logistics (Dovbischuk 2022) and IT (Li et al. 2022), and also prompted a wave of open innovation practices through which inventors widely disseminated their contributions (Brem et al. 2021, Lee and Trimi 2021, Ho 2023). Accounting for innovation shocks is essential. As the COVID-19 surge demonstrated, when a shock occurs, society may be compelled to transform latent capabilities—such as vaccine development or its digital capabilities—into core capabilities (Li et al. 2022, Dovbischuk 2022, Oladapo et al. 2023). Moreover, the timing and adaptability of firms in responding to such shocks –with the consequent propagation of innovative responses via dynamic interactions– are critical to ensuring that newly required capabilities do not merely displace former core capabilities, which, in the aftermath, risk becoming core rigidities (Atanassova and Bednar 2022).

### v2

@@ 1 Introduction
The literature adopts diverse definitions of novelty, innovation, and invention, depending on the degree of originality introduced or the social and economic impact of the advance. In physics and mathematics, innovation processes are modeled as sequences of observed elements: an item is labeled a novelty if it appears for the first time, otherwise it is a repetition (Tria et al. 2014). Classic examples include linguistic innovation, where the introduction of a new word constitutes a novelty. Models such as the Poisson–Dirichlet process (Pitman and Yor 1997) or urn-with-triggering schemes (Tria et al. 2014) formalize these dynamics, and extensions to interacting processes have been developed (Fortini, Petrone and Sporysheva 2018, Iacopini et al. 2020, Aletti, Crimaldi and Ghiglietti 2023, 2025).
In economics, innovation is typically synonymous with patented invention: each patent is novel by definition, but the central question is whether it constitutes a successful innovation. Accordingly, our model treats innovations, novelties, and inventions interchangeably, but focuses on whether a patented item meets a success threshold (with respect to a novelty or citation-based measure). This justifies our choice of interacting Bernoulli processes, which naturally represent success/failure dynamics across categories.
Related approaches exist. Antelman (1972) introduced dependent Bernoulli processes via Dirichlet-beta priors, but restricted to two processes. Katselis, Beck and Srikant (2019) allowed dependence only on contemporaneous states, excluding lagged history. Pandit et al. (2019) modeled dependence on finite lags, requiring sparsity assumptions. By contrast, our model incorporates dependence on the entire past, with parameters that decline over time ($\omega_{kj}/(t+\nu)$), yielding vanishing success probabilities and power-law growth consistent with patent data. Importantly, our theoretical results follow from general asymptotic theorems that encompass and extend prior models, providing a rigorous foundation for studying interacting novelty processes and including earlier contributions as special cases (Tria et al. 2014, Aletti, Crimaldi and Ghiglietti 2023). In particular, we generalize results on the number of distinct observed novelties, situating our model within the broader family of stochastic processes for innovation dynamics.
Outline
@@
The remainder of the paper is structured as follows. Section 2 introduces our novelty measure and defines what constitutes a technological “success” in a technological category. Section 3 presents the mathematical model of interacting Bernoulli processes, establishing its theoretical properties. All the theoretical results are analytically proven in the appendix (supplementary material). Moreover, Section 3 highlights the implications of the model through a simulation exercise, related to the propagation effects of exogenous shocks. In Section 4, we validate the theoretical findings using a comprehensive dataset of patents from the GLOBAL PATSTAT database, and we estimate the structural parameters of the interaction matrix. Statistical inference on the interaction intensity under a mean-field assumption is also performed. Section 5 concludes by discussing implications for innovation policy, capability evolution, and long-run patterns of specialization.
2 Category forward citation index and definition of “success”
@@
As previously noted, in economics the term “innovation” often refers to “patented invention”. Since patents are novel by definition, our focus lies on whether they represent a successful innovation. Thus, we use the terms innovation, novelty, and invention interchangeably, emphasizing their success rather than their degree of novelty.
@@
The number of citations a given patent receives (more briefly, forward citations) reflects the technological importance of the patent for the development of subsequent technologies (see Squicciarini, Dernis and Criscuolo (2013) and reference therein). In other words, forward citations to a certain patent represent the range of later inventions that have benefited from that patent. Therefore, we have decided to use a variant of the forward citation index in Squicciarini, Dernis and Criscuolo (2013) for detecting the patents that could be defined a “success” in one or more categories. However, it is worthwhile to note that our approach is not specific for a certain classification in categories or for the chosen index and so it could be also applied with a different index aimed to evaluate the importance of a patent. [^2] A review of the novelty indices present in the literature and potentially applicable to our context is presented in Appendix A. There we also motivate our preference for the index of Squicciarini, Dernis and Criscuolo (2013) over other measures of patent novelty.
@@
$CIT_{n,h}$= number of patents with publication date in the period $d_{n},d_{n}+365\times T]$ and belonging to category $h$ and citing $n$.
@@
Note that the number of forward citations a patent gets can differ greatly depending on the publication year and the category it comes from. This is why we decided to share the approach adopted in Squicciarini, Dernis and Criscuolo (2013) and normalize this number with respect to the cohort of the patent considered. In other words, the index we use is a disaggregation of the forward citation index, say $I_{S}(n)$, in Squicciarini, Dernis and Criscuolo (2013): indeed, we have
@@
Remark 1 (Invariance property of the index).
@@
According to the above definition, if we denote by $X_{n,h}$ the random variable that is equal to $1$ if patent $n$ is a success for category $h$ and equal to 0 otherwise, we obtain a finite system of Bernoulli processes, that is $\{X^{(h)}=(X_{n,h})_{n}\,:\,h\in{\mathcal{H}}\}$. These processes are not independent, but they interact. This interaction is crucial to consider, as numerous studies have highlighted that –within innovation ecosystems– innovation processes do not evolve in isolation but often influence one another through knowledge spillovers, recombination, and co-dependence, shaping both innovation outcomes and strategic investment decisions across domains (Adner and Feiler 2019, Bondarev and Krysiak 2021, Kim and Magee 2017).
@@
3 Mathematical model and theoretical results: Interacting Bernoulli processes
@@
$P_{t,h}$ $=P(X_{t+1,h}=1\,|\,\text{past until time-step $t$})$ (1)
@@
where $\theta_{h}>0$, $c\geq\theta_{h}$ are parameters that tune the initial condition in the model dynamics (for details, see (4) in Appendix B) and $S_{t,j}=\sum_{n=1}^{t}X_{n,j}$ is the number of successes until time-step $t$ for process $j$. Moreover, the parameters $\gamma_{j,h}$ play the fundamental role in the dynamics of the system (see again (4)). We refer to the matrix $\Gamma=(\gamma_{j,h})_{j,h}$ as the interaction matrix and we assume the following conditions:
@@
    $\Gamma$ is non-negative and such that $\mathbf{1}^{\top}\Gamma\leq\mathbf{1}^{\top}$ (where ${\mathbf{1}}=(1,\ldots,1)^{\top}$), i.e. $\sum_{j\in{\mathcal{H}}}\gamma_{j,h}\leq 1$ for each $h\in{\mathcal{H}}$ so that we have $P_{t,h}\in(0,1)$;
@@
    $\Gamma$ is irreducible, that is the graph with the Bernoulli processes as nodes and with $\Gamma$ as the (weighted) adjacency matrix is strongly connected. (In Appendix C we will discuss the case of a reducible matrix $\Gamma$. [^3] This extension may be particularly relevant in contexts where the USPTO –or other patent offices—modify the structure of technological categories through reclassification (Lafond and Kim 2019, Chae and Gim 2019) thus changing the landscape of interaction among the latter.)
@@
In our context, given $t$ patents, the probability $P_{t,h}$ that the future patent $t+1$ will be a success in category $h$ can have an increasing dependence on the number of past successes in any category (self- and cross-category reinforcement) with the parameter $\gamma_{j,h}$ ruling the strength of this (direct) dependence and the assumed irreducibility of the matrix $\Gamma$ generates a global interaction among all the categories (according to the literature on knowledge spillovers among patents Jaffe, Trajtenberg and Fogarty (2000), Thompson and Fox-Kean (2005)).
@@
  • in the case $\gamma^{*}<1$ (that is when ${\mathbf{1}}^{\top}\Gamma\neq{\mathbf{1}}^{\top}$) the probability of observing a success in process $h$ converges almost surely toward zero at the same rate $1/t^{(1-\gamma^{*})}$ for every process;
  • in the case $\gamma^{*}=1$ (that is when ${\mathbf{1}}^{\top}\Gamma={\mathbf{1}}^{\top}$) the probability of observing a success in process $h$ converges almost surely to a finite strictly positive random limit, which is the same for each process (note that in this case $\mathbf{u}$} is, up to a multiplicative non-zero constant, equal to the vector $\mathbf{1}$}); in other words, at the steady state, the probability of having a success in a process is the same for each process of the system.
@@
Moreover, we recall that, according to the Perron-Frobenius theory, the components of the vector $\mathbf{u}$} are all different from zero and with the same sign and, in graph theory, they correspond to the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$). The eigenvector centrality is a measure of the importance of a node in a graph with respect to its out-links (if it is computed for the adjacency matrix) or its in-links (if it is computed for the transpose of the adjacency matrix). A high eigenvector centrality score means that the node points to, or respectively is pointed to by, many nodes with high scores. In the above theorems $\mathbf{u}$} is the vector of the relative eigenvector centrality scores with respect to $\Gamma^{\top}$ and, hence, by (1), a high value of $u_{h}$ means that the probability of having a success for process $h$ depends on the number of past successes observed in many processes who themselves have high scores.
@@
In short, a balanced innovation ecosystem ($\gamma^{*}=1$) supports ongoing, shared uncertainty and correlated success across technologies, which characterizes connected innovation ecosystems (Jacobides, Cennamo and Gawer 2024). In contrast, a not-balanced one ($\gamma^{*}<1$) leads to asymmetric, decoupled innovation dynamics over time. In other words, in well-connected systems, sustained uncertainty and interdependence across technological domains facilitate diverse learning paths and capability recombination. This interdependence fosters mutual reinforcement across domains, enabling firms to accumulate diverse, complementary capabilities. Conversely, in fragmented systems, limited spillovers may rigidify specialization patterns, potentially turning capabilities into rigidities over time.
@@
It is interesting to note that, in the innovation framework, Theorem 3.2 highlights how the long-run proportions of successful patents in each category reflect the connectedness and influence of that category within the broader innovation network. While this relationship is formally established in the present work, related insights have been suggested in the literature without formal proof. For example, Pichler, Lafond and Farmer (2020) show that the innovation rate of a technological domain is influenced by the innovation rates of the domains it depends on. Similarly, Sampat and Ziedonis (2005), Zhang et al. (2025) emphasize the importance of the quality of a patent’s prior connections as well as their technological domain. In an effort to further deepen this latter aspect, we developed Theorem 3.3. As explained in the following, this theorem examines not only how many successful patents there are in each target category $h$, but also traces the source category $k$ to which each successful patent originally belonged. Namely, in our particular context, we can enrich the model assuming that the number $S_{t,k,h}$ of successes until time-step $t$ in category (process) $h$ coming from belonging category $k$ is of the form $S_{t,k,h}=\sum_{n=1}^{t}X_{n,h}Y_{n,k}$, where $Y_{n,k}$ takes value $1$ if the category of patent $n$ is $k$ and 0 otherwise and
@@
    ${\mathbf{Y}}_{n}=(Y_{n,k})^{\top}_{k\in{\mathcal{H}}}$ is independent of ${\mathbf{X}}_{n}=(X_{n,h})^{\top}_{h\in{\mathcal{H}}}$ and of all the past until time-step n1n-1 with $P(Y_{n,k}=1)=\pi_{k}\in(0,1)$ (where $\sum_{k\in{\mathcal{H}}}\pi_{k}=1$).
@@
To summarize, Theorem 3.3 provides a formal representation of how the structure of innovation inputs shapes the distribution of successful outcomes across technological domains. Specifically, it shows that, asymptotically, the share of successful innovations in a target category mirrors the origin distribution of patent activity. This result offers entrepreneurs a valuable mechanism for understanding how ecosystems embed and propagate specialization trajectories, which is a critical insight, given that the transformation of skills into core capabilities depends fundamentally on how technological domains interact and reinforce one another over time (Teece 2009, Jacobides, Cennamo and Gawer 2018, Kim and Magee 2017).
3.1 Model implication: simulations with a shock
Figure 1: Linear behavior of the successes counter $S_{t,1}$ (red line) and $S_{t,2}$ (blue line) along $t$, in $\log_{10}-\log_{10}$ scale, in two different scenarios with a positive shock on process $1$. The two scenarios are distinguished by the type of interaction, i.e. by the form of the interaction matrix $\Gamma$ used in the simulations. The shock on process 11 occurs at time $t_{shock}=10^{4}$ and the system is observed until time-step $t=10^{7}$. We can observe the sudden rise of $S_{t,1}$ (red line) caused by the shock and how the positive effect of the shock on process $1$ propagates to process 22 (consequent progressive rise of the blue line). We can also note that the slope of the two lines and the distance between their intercepts are the same before and after the shock, because the interaction matrix $\Gamma$ does not change: indeed, the slope is equal to $\gamma^{*}$ and the distance between the intercepts corresponds to the quantity $\log_{10}(u_{1}/u_{2})=\log_{10}(u_{1})-\log_{10}(u_{2})$.
In order to highlight what is the implication of the fact that a system follows the provided model, we here show how the effect of a shock on a process propagates to the other ones by means of the interaction in their dynamics. Having a theoretical model that accounts for the propagation of innovation shocks is essential, as it is a phenomenon observed in real world  [^4] A vivid example of this phenomenon is the surge in pharmaceutical innovation during the COVID-19 pandemic, which triggered subsequent innovations in other fields such as logistics (Dovbischuk 2022) and IT (Li et al. 2022), and also prompted a wave of open innovation practices through which inventors widely disseminated their contributions (Brem, Viardot and Nylund 2021, Lee and Trimi 2021, Ho 2023)..